OpenStudy (redneckoutlaw):

What is the simplified form of x squared plus 6 times x plus 8 all over 4 minus x squared.? a. negative times the quantity x plus 2 all over 4 plus x. b. negative times the quantity x minus 2 all over 4 minus x. c. x plus 4 all over 4 minus x. d. negative times the quantity x plus 4 all over x minus 2.

OpenStudy (anonymous):

\[x ^{2}+6x + 8 / 4-x ^{2}\]

OpenStudy (anonymous):

Is that right, for the non-simplified version?

OpenStudy (redneckoutlaw):

Yes

OpenStudy (anonymous):

We just have to factor the quadratic in the numerator, and the difference of squares in the denominator.

OpenStudy (anonymous):

\[\frac{ x^2 +6x + 8 }{ 4 - x^2 }\] Factoring out :\[x^2 + 6x + 8 => x^2 + 4x + 2x + 8 => x(x+4) +2(x+ 4) => x(x+4)(x+2)\] \[\frac{ (x+4)(x+2) }{ (x+4) } = (x+2)\]

OpenStudy (anonymous):

So, x^2+6x+8 becomes (x+4)(x+2) and 4-x^2 becomes (2-x)(2+x)

OpenStudy (anonymous):

@RedNeckOutLaw

OpenStudy (anonymous):

Then we cancel out the common factor of (x+4), leaving us with (x+2)

OpenStudy (redneckoutlaw):

Cos, wouldnt it be (x+2) (x+2)

OpenStudy (anonymous):

No. A difference of squares will always yield a + and - sign.

OpenStudy (anonymous):

4-x^2 = (x+2)(x-2) x^2-9 = (x+3)(x-3) etc.

OpenStudy (redneckoutlaw):

Ohhh, i understand where yoru comignfrom

OpenStudy (anonymous):

Remember, when it's 2 squares, like x^2 and a 3, 9, 16, 25, 36, etc. and you are taking the difference, Example, x^2-9 or 9-x^2 x^2-25 or 25-x^2 It has to be be (x+3)(x-3) or (3+x)(3-x) (x+5)(x-5) or (5+x)(5-x) Where it has to be a + times a -, to give a negative.

OpenStudy (anonymous):

=)

OpenStudy (redneckoutlaw):

So fir the first example it can either way. so f its x^2-25 or 25-x^2. its (x+5)(x-5) or (5+x)(5-x) its because 5x5 is 25

OpenStudy (redneckoutlaw):

@Cosmichaotic

OpenStudy (anonymous):

Exactly, yes.

OpenStudy (redneckoutlaw):

Ok, so whats the next thing we do?

OpenStudy (anonymous):

Well, in your problem, it becomes